# Remasking Discrete Diffusion Models with Inference-Time Scaling

### v2

@@ Appendix C ReMDM Sampler Algorithms
        $\sigma_{t}^{(\ell)}=\eta_{rescale}\cdot\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$ for all =$\ell=\{1,\ldots,L\}$.
@@
=$=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t}^{(\ell)})\mathbf{x}_{\theta}^{(\ell)}+\sigma_{t}^{(\ell)}\bm{m}),&\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t}^{(\ell)})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}_{\theta}^{(\ell)}+\frac{1-\alpha_{s}-\sigma_{t}^{(\ell)}\alpha_{t}}{1-\alpha_{t}}\bm{m}),&\mathbf{z}_{t}=\bm{m}\end{cases}$
@@ D.1 OpenWebText
        $\sigma_{t}^{(\ell)}=\eta_{rescale}\cdot\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$ for all =$\ell=\{1,\ldots,L\}$.
@@
=$=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t}^{(\ell)})\mathbf{x}_{\theta}^{(\ell)}+\sigma_{t}^{(\ell)}\bm{m}),&\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t}^{(\ell)})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}_{\theta}^{(\ell)}+\frac{1-\alpha_{s}-\sigma_{t}^{(\ell)}\alpha_{t}}{1-\alpha_{t}}\bm{m}),&\mathbf{z}_{t}=\bm{m}\end{cases}$
@@ D.2 ImageNet
In this experiment, we reuse the pre-trained MaskGiT model for all samplers [7]. The MaskGiT architecture is a Transformer model [53] that consists of 24 layers, 8 attention heads, 768 embedding dimensions, and 3,072 hidden dimensions. Similarly to MDLM, this model implements the carry-over unmasked tokens property in the parameterization of $\mathbf{x}_{\theta}$. The full model architecture and training setup details are available in Chang et al. [7].
@@ D.4 LLaDA
We reuse the open-source LLaDA 8B Instruct model from [28] in this experiment. Since LLaDA utilizes MaskGiT-style samplers, we modify ReMDM-loop to similar samplers with confidence heuristics. For a maximum mask length equal to 32, we first use the LLaDA sampler to generate 28 tokens. Then we step into the ReMDM-loop phase, where we remask a fixed number (1 for Countdown and 4 for TruthfulQA) of tokens with the smallest decoding confidence and unmask the same number of tokens each time step. The ReMDM-loop phase lasts 32 steps, and after the loop phase, we use the LLaDA sampler to fill in the remaining gaps. For DFM, since it does not support looping (see Proposition 4.4), we remove the loop phase and let it remask one token and unmask two tokens each time step instead. Different from LLaDA, which utilizes greedy sampling, we enable random sampling to compute error bars.
We use the widely used evaluation harness [12] as the evaluation framework, and for both Countdown and TruthfulQA, we use a maximum length of 32 tokens and no semi-autoregressive sampling introduced in [28], that is, the block size equals 32 due to relatively short answer lengths. For Countdown, we use 4-shot evaluation, and for TruthfulQA, we use 6-shot evaluation (see Appendix F.3 for the chat templates we use). For Countdown, we evaluate on 256 synthetically generated countdown questions. We repeat the sampling process for each question 10 times and calculate pass@1 as $\frac{c}{10}$ where $c$ is the number of correct trials. After this, we report the average pass@1 for the 256 questions. For TruthfulQA, we let the model answer 817 questions on factual knowledge and calculate the ROUGE scores between the model and correct answers, as well as the model and plausibly wrong answers. The difference between the right and wrong ROUGE scores is reported. Due to the low computational resource demand for the two tasks (one run takes around 30 minutes using one Nvidia A100 GPU), for both tasks, we repeat the experiments 20 times with 20 different random seeds (0-19) and report the mean values and 95% confidence intervals.
@@ Appendix E Additional Experimental Results
In Table 4, we report QM9 dataset validation set perplexities for models trained with either the MDLM NELBO from (3) or the ReMDM NELBO from (3.2), with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$, for some $\eta_{cap}\in(0,1]$. We follow the same model architecture and training settings defined in Schiff et al. [43] for this dataset. Overall, we find that validation perplexities for this dataset are fairly consistent across models trained with either objective.
Table 4: Training with ReMDM and MDLM NELBO objectives leads to similar validation set perplexity values (for the QM9 dataset). In the top row, the model is trained using the continuous time formulation of the MDLM objective from Sahoo et al. [38]. In the bottom two rows, models are trained with discrete-time objectives. The first column, $\sigma_{t}=0$, corresponds to the MDLM objective from (3) and the columns with varying $\eta_{cap}$ correspond to training with the ReMDM objective from (3.2) with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$. We find that results are comparable across training settings.

### v3

@@ Appendix C ReMDM Sampler Algorithms
   $\sigma_{t}^{(\ell)}=\eta_{rescale}\cdot\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$ for all $\ell=\{1,\ldots,L\}$.
@@
$=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t}^{(\ell)})\mathbf{x}_{\theta}^{(\ell)}+\sigma_{t}^{(\ell)}\bm{m}),&\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t}^{(\ell)})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}_{\theta}^{(\ell)}+\frac{1-\alpha_{s}-\sigma_{t}^{(\ell)}\alpha_{t}}{1-\alpha_{t}}\bm{m}),&\mathbf{z}_{t}=\bm{m}\end{cases}$
@@ D.1 OpenWebText
   $\sigma_{t}^{(\ell)}=\eta_{rescale}\cdot\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$ for all $\ell=\{1,\ldots,L\}$.
@@
$=\begin{cases}\mathrm{Cat}(\mathbf{z}_{s};(1-\sigma_{t}^{(\ell)})\mathbf{x}_{\theta}^{(\ell)}+\sigma_{t}^{(\ell)}\bm{m}),&\mathbf{z}_{t}\neq\bm{m}\\ \mathrm{Cat}(\mathbf{z}_{s};\frac{\alpha_{s}-(1-\sigma_{t}^{(\ell)})\alpha_{t}}{1-\alpha_{t}}\mathbf{x}_{\theta}^{(\ell)}+\frac{1-\alpha_{s}-\sigma_{t}^{(\ell)}\alpha_{t}}{1-\alpha_{t}}\bm{m}),&\mathbf{z}_{t}=\bm{m}\end{cases}$
@@ D.2 ImageNet
In this experiment, we reuse the pre-trained MaskGiT model for all samplers [7]. The MaskGiT architecture is a Transformer model [56] that consists of 24 layers, 8 attention heads, 768 embedding dimensions, and 3,072 hidden dimensions. Similar to MDLM, this model implements the carry-over unmasked tokens property in the parameterization of $\mathbf{x}_{\theta}$. The full model architecture and training setup details are available in Chang et al. [7].
@@ D.4 LLaDA
We reuse the open-source LLaDA 8B Instruct model from Nie et al. [31] in this experiment. Since LLaDA utilizes MaskGiT-style samplers, we modify ReMDM-loop to similar samplers with confidence heuristics. For a maximum mask length equal to 32, we first use the LLaDA sampler to generate 28 tokens. Then we step into the ReMDM-loop phase, where we remask a fixed number (1 for Countdown and 4 for TruthfulQA) of tokens with the smallest decoding confidence and unmask the same number of tokens each time step. The ReMDM-loop phase lasts 32 steps, and after the loop phase, we use the LLaDA sampler to fill in the remaining gaps. For DFM, since it does not support looping (see Proposition 4.4), we remove the loop phase and let it remask one token and unmask two tokens each time step instead. Different from LLaDA, which utilizes greedy sampling, we enable random sampling to compute error bars.
We use the widely adopted evaluation harness from Gao et al. [12], and for both Countdown and TruthfulQA, we use a maximum length of 32 tokens without semi-autoregressive sampling introduced in [31], that is, the block size equals 32 due to relatively short answer lengths. For Countdown, we use 4-shot evaluation, and for TruthfulQA, we use 6-shot evaluation (see Appendix F.3 for the chat templates we use). For Countdown, we evaluate on 256 synthetically generated countdown questions. We repeat the sampling process for each question 10 times and calculate pass@1 as $\frac{c}{10}$ where $c$ is the number of correct trials. We then report the average pass@1 for the 256 questions. For TruthfulQA, we let the model answer 817 questions on factual knowledge and calculate the ROUGE scores between the model’s answers and correct answers, as well as the model and plausibly wrong answers. The difference between the correct and incorrect ROUGE scores is reported. For both tasks, we repeat the experiments 20 times with 20 different random seeds (0-19) and report the mean values and 95% confidence intervals.
@@ Appendix E Additional Experimental Results
In Table 4, we report QM9 dataset validation set perplexities for models trained with either the MDLM NELBO from (3) or the ReMDM NELBO from (9), with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$, for some $\eta_{cap}\in(0,1]$. We follow the same model architecture and training settings defined in Schiff et al. [46] for this dataset. Overall, we find that validation perplexities for this dataset are fairly consistent across models trained with either objective.
Table 4: Training with ReMDM and MDLM NELBO objectives leads to similar validation set perplexity values (for the QM9 dataset). In the top row, the model is trained using the continuous time formulation of the MDLM objective from Sahoo et al. [41]. In the bottom two rows, models are trained with discrete-time objectives. The first column, $\sigma_{t}=0$, corresponds to the MDLM objective from (3) and the columns with varying $\eta_{cap}$ correspond to training with the ReMDM objective from (9) with $\sigma_{t}=\min\{\eta_{cap},(1-\alpha_{s})/\alpha_{t}\}$. We find that results are comparable across training settings.